OpenStudy (anonymous):

A tandem (two-person) bicycle team must overcome a force of 34 lb to maintain a speed of 30 ft/s. Find the power required per rider, assuming they each contribute equally. Express your answer in horsepower.

OpenStudy (anonymous):

I do not know these conversions but I can tell you how to do the problem..

OpenStudy (anonymous):

Power = Force x velocity Since, the power here is for two persons, divide the power by 2 to get individual power as it is said that they contribute equally....then convert your answer into horse power

OpenStudy (anonymous):

okay thanks!! :) can i ask one more question?

OpenStudy (anonymous):

yea..post it

OpenStudy (anonymous):

A pump is required to lift 200 L of water per minute from a wall 10 m deep and eject it with a speed of 20 m/s. (a) How much work is done per minute in lifting the water? (b) How much in giving it kinetic energy? (c) What horsepower engine is needed if it is 80% efficient?

OpenStudy (anonymous):

density of water is 1 g/ml Mass of water(in grams) = Volume of water(in ml) x 1(density) a) work done = mgh m : mass of water g : acceleration due to gravity = 9.8 m/s^2 or 10 m/s^2 h : height or depth = 10 m b) kinetic energy = (1/2)mv^2 m : mass of water v : velocity of ejection = 20 m/s c) If the engine is 80% efficient, then 80% of its capacity should equal to the work obtained in part a. Let the engine's actual work be W (80/100) x W = work calculated in part a.